# Navigate the Unknown: Enhancing LLM Reasoning with Intrinsic Motivation Guided Exploration

### v5

@@ Abstract
Reinforcement Learning (RL) has emerged as a pivotal method for improving the reasoning capabilities of Large Language Models (LLMs). However, prevalent RL approaches such as Proximal Policy Optimization (PPO) and Group-Regularized Policy Optimization (GRPO) face critical limitations due to their reliance on sparse outcome-based rewards and inadequate mechanisms for incentivizing exploration. These limitations result in inefficient guidance for reasoning. Specifically, sparse rewards fail to deliver sufficient feedback, particularly for challenging problems. Furthermore, such rewards induce systematic biases that prioritize exploitation of familiar trajectories over novel solution discovery. These shortcomings critically hinder performance in complex reasoning tasks, which inherently demand iterative refinement across intermediate steps. To address these challenges, we propose an Intrinsic Motivation guidEd exploratioN meThOd foR LLM Reasoning (i-MENTOR), a method designed to deliver dense rewards and amplify exploration in the RL-based paradigm. i-MENTOR introduces three innovations: trajectory-aware exploration rewards that mitigate bias in token-level strategies while maintaining computational efficiency; error-conditioned reward allocation to ensure efficient exploration on challenging samples while intrinsically stabilizing training; and advantage-preserving integration that maintains advantage distribution integrity while incorporating exploratory guidance. Experiments across 4 public datasets demonstrate i-MENTOR’s effectiveness, achieving a 22.23% improvement on AIME 2024. The source code is available for reference  [^1] https://github.com/Gaojingtong/iMENTOR.
@@ 1 Introduction
Reinforcement learning (RL) (Arulkumaran et al., 2017; Shakya et al., 2023; Ladosz et al., 2022) has become essential for training Large Language Models (LLMs) (Hadi et al., 2023; Min et al., 2023), evolving from Proximal Policy Optimization (PPO) (Schulman et al., 2017) to Group-Regularized Policy Optimization (GRPO) (Shao et al., 2024). Frameworks like DeepSeek (Guo et al., 2025; Liu et al., 2024a; Shao et al., 2024) showcase State-Of-The-Art (SOTA) implementations that optimize multi-step reasoning using trajectory-level advantage estimation. RL uniquely transforms sparse rewards (Gao et al., 2024) into learning gradients, enabling coherent Chain-of-Thought (CoT) (Wei et al., 2022; Plaat et al., 2024) reasoning. By refining action distributions through feedback, RL bridges the gap between linguistic competence and goal-directed problem-solving in LLMs.
@@
While contemporary RL methods like PPO and GRPO demonstrate progress through outcome-based rewards (Uesato et al., 2022; Lightman et al., 2023; Wang et al., 2024a), two fundamental limitations persist for reasoning tasks. First, the reliance on static reward functions creates a sparse learning signal (Bayat et al., 2025; Cao et al., 2024) that fails to guide intermediate reasoning steps, forcing models to navigate vast action spaces with terminal outcome feedback alone. This is particularly evident on difficult samples and datasets (Liu et al., 2024b) where outcome-based rewards are more likely to be zero. Second, despite GRPO’s group-wise sampling strategy, the inherent reward structure disincentivizes genuine exploration – trajectories yielding identical final outcomes receive equivalent advantage estimates regardless of different reasoning CoTs, effectively penalizing computationally intensive sampling efforts (Yu et al., 2025; Liu et al., 2025). This creates a paradoxical scenario where models optimize for reward exploitation at the expense of systematic exploration, particularly detrimental in difficult reasoning tasks that require complicated reasoning processes.
Traditional exploration methods (e.g., RND(Burda et al., 2018), ICM(Pathak et al., 2017), Count-Based Exploration (Ostrovski et al., 2017)) encourage agents to explore novel or under-visited states via intrinsic rewards, inspired by cognitive theories of curiosity. While promising for guiding LLM exploration, these methods face challenges in LLM reasoning tasks due to: 1) Dynamic episodic length and computational overload arise due to dynamic CoT length and token-level exploration rewards (Burda et al., 2018). Directly using each reasoning step as a sample for traditional exploration methods will result in the exploration reward of the long sequence being higher than that of the short sequence, inducing the model to explore samples with longer sequences. Moreover, the per-token computation for lengthy LLM outputs incurs significant computational overhead. 2) Large action space poses a challenge due to the exponential growth in possible reasoning paths for LLMs (Zhou et al., 2024; Wen et al., 2024). Naive exploration strategies may incur prohibitive computational costs, negligible outcomes, and even training instability. 3) Exploration reward integration conflicts with outcome rewards in methods such as GRPO (Shao et al., 2024). The reward normalization in GRPO may invert advantages from positive to negative when exploration bonuses are included, disproportionately penalizing low-exploration samples instead of incentivizing high-value exploration. This introduces destabilizing noise to policy optimization and undermines outcome reward efficacy.
In this paper, we propose an Intrinsic Motivation guidEd exploratioN meThOd foR LLM Reasoning (i-MENTOR), which mitigates reward sparsity, enables smart explorations in LLM training, and effectively resolves the above challenges: 1) Trajectory-aware exploration rewards operate at the sequence level to eliminate computational overload and sequence-length bias caused by dynamic episodic length. By employing two lightweight trajectory-aware networks, this component efficiently captures reasoning sequence uniqueness while maintaining computational traceability. 2) Error-conditioned reward allocation selectively applies exploration incentives exclusively to incorrect reasoning trajectories and improves exploration efficiency while maintaining a stable training process. This enables efficient exploration of challenging samples while intrinsically stabilizing training. 3) Advantage-preserving integration introduces exploration incentives after advantage computation. This approach resolves inherent conflicts between exploration rewards and outcome-based rewards. By preserving the statistical integrity of outcome-driven advantage distributions, it integrates exploratory guidance to enable effective coordination between these objectives. By unifying the above mechanisms, i-MENTOR generates dense, stable, and computationally efficient exploration rewards that seamlessly integrate into RL-based LLM training pipelines. This enables enhanced reasoning capabilities through balanced exploration-exploitation dynamics, overcoming the limitations of conventional exploration rewards in LLM-based sequential decision-making contexts.
Our contributions are summarized as follows: (1) We propose i-MENTOR, a novel method that systematically incorporates dense, stable, and computationally efficient exploration rewards to enhance LLM reasoning. Designed for seamless integration with established RL algorithms, i-MENTOR leverages exploration rewards to improve learning efficiency while maintaining stable policy updates throughout the training process. (2)We present three advances: a) Trajectory-aware rewards generate exploration rewards while maintaining efficiency; b) Error-conditioned reward allocation ensures efficient exploration on challenging samples while intrinsically stabilizing training; c) Advantage-preserving integration injects exploration rewards after advantage computation to preserve policy gradient fidelity. Together, i-MENTOR systematically mitigates reward sparsity and enhances LLM reasoning through coordinated explorations. (3) Experiments on four public datasets and two base LLMs (i.e., Qwen2.5-3B (Team, 2024) and deepseek-7b (DeepSeek-AI, 2025)) demonstrate i-MENTOR’s consistent effectiveness across various RL methods and base models, significantly enhancing the reasoning capabilities of LLMs. (4) Our experiments reveal an interesting pattern: i-MENTOR’s improvements are more pronounced on difficult datasets. The results suggest that i-MENTOR’s dense exploration rewards for each response sequence specifically help models overcome learning barriers in challenging samples. This phenomenon represents a further contribution of our study: it originates from Reinforcement Learning with Verifiable Rewards (RLVR)’s core distinction from traditional RL–leveraging a pre-trained LLM instead of a randomly initialized model. Consequently, exploration yields significantly greater benefits when addressing complex problems.
@@ 2 Method
In this section, we first briefly introduce the current SOTA RL training methods for LLM, i.e., PPO and GRPO, then introduce our method, i-MENTOR.
@@ 2.1.1 Proximal Policy Optimization (PPO)
$\mathcal{J}_{\mathrm{PPO}}(\theta)=\mathbb{E}_{(q,a)\sim\mathcal{D},o_{\leq t}\sim\pi_{\theta_{\text{old }}}(\cdot\mid q)}\left[\min\left(w_{t}(\theta)\hat{A}_{t},\operatorname{clip}\left(w_{t}(\theta),1-\varepsilon,1+\varepsilon\right)\hat{A}_{t}\right)\right]$ (1)
where $w_{t}(\theta)=\frac{\pi_{\theta}\left(o_{t}\mid q,o_{<t}\right)}{\pi_{\theta_{\text{old }}}\left(o_{t}\mid q,o_{<t}\right)}$. Here, $(q,a)$ denotes question-answer pairs from dataset $\mathcal{D}$. $o_{t}$ and $o_{<t}$ are generated responses end at token position $t$ and t1t-1. The current and previous policies are parameterized by $\pi_{\theta}$ and $\pi_{\theta_{\text{old}}}$ respectively. The advantage estimator $\hat{A}_{t}$ employs Generalized Advantage Estimation (GAE) (Schulman et al., 2015) with outcome-based reward function $R$ and value function $V$ for advantage estimation, while $\varepsilon$ controls the clipping range. By treating question-response sequences ending at each token position of $o_{t}$ as distinct states and constraining policy updates within a trust region, PPO stabilizes RL training for LLMs. However, the joint optimization of policy and value functions introduces computational overhead, limiting training efficiency.
2.1.2 Group-Regularized Policy Optimization (GRPO)
@@
$\mathcal{J}_{\mathrm{GRPO}}(\theta)$ $=\mathbb{E}_{(q,a)\sim\mathcal{D},\left\{o_{i}\right\}_{i=1}^{G}\sim\pi_{\theta_{\text{old }}(\cdot\mid q)}}$ (2)
${\left[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{\left|o_{i}\right|}\sum_{t=1}^{\left|o_{i}\right|}\left(\min\left(w_{i,t}(\theta)\hat{A}_{i},\operatorname{clip}\left(w_{i,t}(\theta),1-\varepsilon,1+\varepsilon\right)\hat{A}_{i}\right)-\beta D_{\mathrm{KL}}\left(\pi_{\theta}\|\pi_{\mathrm{ref}}\right)\right)\right].}$
where $w_{i,t}(\theta)=\frac{\pi_{\theta}\left(o_{i,t}\mid q,o_{i,<t}\right)}{\pi_{\theta_{\text{old }}}\left(o_{i,t}\mid q,o_{i,<t}\right)}$, $\hat{A}_{i}=\frac{R_{i}-\operatorname{mean}\left(\left\{R_{i}\right\}_{i=1}^{G}\right)}{\operatorname{std}\left(\left\{R_{i}\right\}_{i=1}^{G}\right)}$, $G$ responses $\{o_{i}\}_{i=1}^{G}$ are generated per input, $\pi_{\mathrm{ref}}$ denotes the reference policy, and $\beta$ controls KL penalty strength. GRPO replaces value function estimation with group-wise normalization on outcome-based rewards, improving training efficiency and effects through reduced computational complexity and group-wise sampling process.
Figure 1: i-MENTOR. Values with green/yellow boxes denote higher outcome-based rewards/ advantages and exploration rewards; red/blue denote lower values. Black boxes mark the excluded values. $\hat{A}_{old}$ and $\hat{A}_{new}$ denote advantages derived from outcome-based rewards and i-MENTOR.
2.2 i-MENTOR
Despite advancements, PPO and GRPO face two key limitations: (1) Sparse reward signals that provide limited training guidance, and (2) weak exploration mechanisms during policy updates. These issues hinder learning efficiency and final optimization. To address these challenges, we propose an Intrinsic Motivation guidEd exploratioN meThOd foR LLM Reasoning (i-MENTOR), which introduces structured exploration to enhance LLM reasoning and can be easily implemented in SOTA RL methods like PPO and GRPO. As shown in Figure 1, vanilla RL methods with outcome-based rewards focus only on correct trajectories, ignoring under-visited ones and causing LLMs to converge suboptimally. i-MENTOR instead rewards incorrect under-visited trajectories, preventing the model from getting stuck via intrinsic motivation guided exploration.
A Case Study is presented in Figure 2 to illustrate the significance of i-MENTOR. Specifically, vanilla GRPO-trained LLM exhibits logical errors (e.g., missing numbers, redundancy in ①), repetitive reasoning patterns (②) due to intensive exploitations without explorations, and failures to solve tasks (③) or apply critical operations like multiplication (④). With the guidance of the exploration rewards from i-MENTOR, LLM trained with i-MENTOR-GRPO reduces logical errors, diversifies reasoning paths (④), and results in a correct answer (③), demonstrating improved reasoning ability. Additionally, as shown in Figure 3(a), i-MENTOR increases the average response length during training, stabilizing at a higher level than vanilla RL methods. This suggests that i-MENTOR ’s exploration strategies enable LLMs to learn more complex reasoning processes with longer CoTs through considering diverse reasoning paths in responses, leading to performance gains. In the following subsections, we will systematically introduce i-MENTOR through detailing its key components.
@@
(a) Response length.
(b) RND input statistics.
Figure 3: Training statistics on Countdown-4 (Gandhi et al., 2024).
@@ 2.2.1 Trajectory-aware Exploration Rewards
We propose i-MENTOR to provide intrinsic exploration rewards for LLM reasoning based on the Random Network Distillation (RND) framework (Burda et al., 2018) as it is computationally efficient and does not need prior knowledge.
RND Network: The original RND formulation uses a randomly initialized frozen target network $f_{\theta_{T}}$ and a trainable predictor $f_{\theta_{P}}$ (same architecture as $f_{\theta_{T}}$), minimizing prediction error on visited training states:
$\mathcal{L}(\boldsymbol{s}_{t})=\|f_{\theta_{P}}(\boldsymbol{s}_{t})-f_{\theta_{T}}(\boldsymbol{s}_{t})\|_{2}^{2}$ (3)
where $\boldsymbol{s}_{t}$ denotes the state at step $t$ ($\boldsymbol{s}_{t}=[q,o_{t}]$, part of the question-answer sequence for LLM). The predictor updates concurrently with the main RL algorithm. With the same architecture as $f_{\theta_{T}}$, $f_{\theta_{P}}$ could converges toward $f_{\theta_{T}}$ independently of the RL training models. This enables $\mathcal{L}(\boldsymbol{s}_{t})$ to serve as an intrinsic reward for exploring novel, high-uncertainty states: a high $\mathcal{L}(\boldsymbol{s}_{t})$ indicates poor convergence of $f_{\theta_{P}}$, reflecting insufficient exploration on similar sequences; conversely, a low loss implies strong familiarity. The exploration reward is consequently defined as:
$R^{rnd}(\boldsymbol{s}_{t})=\frac{\|f_{\theta_{P}}(\boldsymbol{s}_{t})-f_{\theta_{T}}(\boldsymbol{s}_{t})\|_{2}^{2}}{\operatorname{std}\left(\|f_{\theta_{P}}(\boldsymbol{s}_{t})-f_{\theta_{T}}(\boldsymbol{s}_{t})\|_{2}^{2}\right)}.$ (4)
@@
The denominator uses a moving average of standard deviations during training for a consistent reward scale and temporal stability, but remains fixed during evaluation for determinism.
However, directly applying such a method to LLMs leads to catastrophic performance and efficiency degradation. For LLM reasoning where states correspond to tokens $\boldsymbol{s}_{t}=[q,o_{t}]$, direct RND application introduces two challenges: (1) Dynamic episodic length: Sequences $(q,o)$ with dynamic CoT length in response $o$ generate inputs $\{\boldsymbol{s}_{t}\}_{t=1}^{|o|}$ with lengths $|q|+1$ to $|q|+|o|$ for RND update. This results in systematic under-sampling of longer input sequences, as they only appear in extended responses, while shorter inputs recur across all sequences regardless of response length, as illustrated in Figure 3(b). (2) Computational overload: Processing thousands of tokens per sequence requires excessive RND network updates.
These phenomena are particularly evident in Figure 3(b). In this figure, a total of $253+416+515+549+701=2434$ RND inputs are generated from merely 5 rollout responses of 1 question with length $[253,416,515,549,701]$ in GRPO optimization. Shorter input states dominate these samples, inducing lower exploration rewards from frequent visits. RND requires multiple forward passes per sequence, creating computational overload.
To address these challenges, we propose trajectory-aware exploration rewards operating at the sequence level. Our approach maintains the dual-network architecture but processes complete sequences $(q,o)$ instead of multiple token-level states $\{\boldsymbol{s}_{t}\}_{t=1}^{|o|}$ as inputs for predictor / target networks:
@@
$\left\{\begin{aligned} &\mathcal{L}(q,o)=\|f_{\theta_{P}}(q,o)-f_{\theta_{T}}(q,o)\|_{2}^{2}\\ &R^{\star}_{1}(q,o)=\frac{\|f_{\theta_{P}}(q,o)-f_{\theta_{T}}(q,o)\|_{2}^{2}}{\operatorname{std}\left(\|f_{\theta_{P}}(q,o)-f_{\theta_{T}}(q,o)\|_{2}^{2}\right)}\end{aligned}\right$. (5)
@@
Here $\mathcal{L}(q,o)$ is the loss function for updating $f_{\theta_{P}}$ while $R^{\star}_{1}(q,o)$ provides a single exploration reward per sequence. This design eliminates token-level bias through uniform sequence treatment and reduces computational complexity from $O(|o|)$ to $O(1)$ per training sequence.

### v6

@@ Abstract
Reinforcement Learning (RL) has become a key approach for enhancing the reasoning capabilities of large language models. However, prevalent RL approaches like proximal policy optimization and group relative policy optimization suffer from sparse, outcome-based rewards and weak exploration incentives, limiting their effectiveness. Specifically, sparse rewards offer limited feedback, especially on difficult problems, and introduce biases favoring familiar trajectories over novel reasoning paths. These issues critically undermine performance on complex tasks that inherently require iterative reasoning. To overcome these challenges, we propose Intrinsic MotivAtion Guided exploratIoN for Enhanced reasoning (IMAGINE), which delivers dense rewards and encourages exploration. IMAGINE introduces three innovations: a trajectory-aware exploration reward that reduces token-level bias efficiently; an error-conditioned reward allocation that promotes efficient exploration on hard samples while stabilizing training; and an advantage-preserving integration mechanism that retains distributional integrity during learning. Experiments on four public datasets show that IMAGINE improves performance by 22.23% on AIME 2024. The source code is available to support reproducibility  [^1] https://github.com/Gaojingtong/IMAGINE.
Machine Learning, ICML
@@ 1 Introduction
Reinforcement Learning (RL) (Arulkumaran et al., 2017; Shakya et al., 2023; Ladosz et al., 2022) has become essential for training Large Language Models (LLMs) (Hadi et al., 2023; Min et al., 2023), evolving from Proximal Policy Optimization (PPO) (Schulman et al., 2017) to Group Relative Policy Optimization (GRPO) (Shao et al., 2024). Frameworks like DeepSeek (Guo et al., 2025; Liu et al., 2024a; Shao et al., 2024) showcase State-Of-The-Art (SOTA) methods that optimize multi-step reasoning using trajectory-level advantage estimation. RL uniquely transforms sparse rewards (Gao et al., 2024) into learning gradients, enabling coherent Chain-of-Thought (CoT) reasoning (Wei et al., 2022; Plaat et al., 2024). By refining action distributions through feedback, RL bridges the gap between linguistic competence and goal-directed problem-solving in LLMs.
@@
While contemporary RL methods like PPO and GRPO demonstrate progress through outcome-based rewards (Uesato et al., 2022; Lightman et al., 2023; Wang et al., 2024a), two fundamental limitations persist for reasoning tasks. First, the reliance on static reward functions creates a sparse learning signal (Bayat et al., 2025; Cao et al., 2024) that fails to guide intermediate reasoning steps due to the limited meaningful rewards on hard samples and early training stages, forcing models to navigate vast action spaces with terminal outcome feedback alone (Bougie and Watanabe, 2025; Song et al., 2025; Bamba and others, 2025). This is particularly evident on difficult samples and datasets (Liu et al., 2024b) where outcome-based rewards are more likely to be zero. Second, despite GRPO’s group-wise sampling strategy, the inherent reward structure disincentivizes genuine exploration – trajectories yielding identical final outcomes receive equivalent advantage estimates regardless of different reasoning CoTs, effectively penalizing computationally intensive sampling efforts (Yu et al., 2025; Liu et al., 2025). This creates a paradoxical scenario where models optimize for reward exploitation at the expense of systematic exploration, particularly detrimental in difficult reasoning tasks that require complicated reasoning processes.
Traditional exploration methods (e.g., RND (Burda et al., 2018), ICM (Pathak et al., 2017), Count-Based Exploration (Ostrovski et al., 2017)) promote novelty-seeking via intrinsic rewards, motivated by cognitive theories of curiosity. However, once exploration is moved from embodied control to LLM reasoning, the usual assumptions begin to break: “progress” is judged on complete responses and shrinks with problem difficulty, and the state space is combinatorial. In this setting, directly reusing token-level exploration can unintentionally introduce bias towards lengthy responses due to unbalanced sampling, inflate per-token computation when reasoning becomes long, and spread exploration over superficial next tokens rather than sequence-level semantics (Zhou et al., 2024; Wen et al., 2024).
A second challenge arises when integrating intrinsic rewards with advanced Reinforcement Learning with Verifiable Rewards (RLVR) objectives. Methods such as GRPO (Shao et al., 2024) rely on carefully normalized advantage estimates to ensure stable policy updates. When exploration rewards are mixed into the same update stream too early, they can alter subsequent advantage statistics in undesirable ways: sometimes punish trajectories simply because less “novelty” may reverse the update direction in advantage estimation. Subsequently, instead of serving as an incentive guide through sparse reward spaces, such exploration can become a competing and distorting objective that alters optimization dynamics.
In this paper, we propose an Intrinsic MotivAtion Guided exploratIoN for Enhanced reasoning (IMAGINE), which makes intrinsic motivation compatible with the way reasoning is expressed and optimized in RL-trained LLMs. IMAGINE replaces token-centric novelty with sequence-level signals by using two lightweight trajectory-aware networks to characterize the uniqueness of complete reasoning paths, avoiding both length bias and token-wise overhead. It further focuses exploration where it is most useful by allocating intrinsic incentives only to incorrect trajectories, creating a self-regulating pressure that concentrates on hard cases while remaining stable as the model improves. Finally, IMAGINE composes intrinsic incentives with outcome learning in a gradient-safe manner: exploration bonuses are introduced only after advantage computation, preserving the outcome-driven advantage distribution while still steering the search toward diverse reasoning strategies. Together, these choices produce dense, stable, and computationally efficient exploration rewards that integrate cleanly into RL-based LLM training pipelines.
Our contributions are as follows. We introduce IMAGINE, an exploration method for RL-based LLM reasoning that provides dense intrinsic guidance without sacrificing training stability or computational efficiency. The method couples trajectory-aware exploration modeling with an error-conditioned reward allocation scheme and an advantage-preserving integration strategy, yielding exploration signals that stay aligned with outcome optimization. Empirically, across four public datasets and two base LLMs (Qwen2.5-3B (Team, 2024) and DeepSeek-7B (DeepSeek-AI, 2025)), IMAGINE consistently improves LLM’s reasoning performance under multiple RL algorithms, including PPO (Schulman et al., 2017), GRPO (Shao et al., 2024), and DAPO (Yu et al., 2025). Notably, the gains are strongest on more difficult benchmarks, suggesting that such sequence-level intrinsic rewards are particularly effective at helping pretrained policies overcome the learning barriers that arise in challenging reasoning instances under RLVR fine-tuning.
@@ 2 Method
Here we first introduce the commonly used RL methods for LLM, i.e., PPO and GRPO, then introduce our method IMAGINE.
@@ 2.1 Preliminary
Figure 1: IMAGINE. Values with green/yellow boxes denote higher outcome rewards/advantages and exploration rewards; red/blue denote lower values. Black boxes mark the excluded values. $\hat{A}_{old}$ and $\hat{A}_{new}$ denote advantages derived from outcome rewards and IMAGINE.
@@ 2.1.1 Proximal Policy Optimization (PPO)
$\begin{split}\mathcal{J}_{\mathrm{PPO}}(\theta)=\mathbb{E}_{(q,a)\sim\mathcal{D},o_{\leq t}\sim\pi_{\theta_{\text{old }}}(\cdot\mid q)}[\min(w_{t}(\theta)\hat{A}_{t},\\ \operatorname{clip}(w_{t}(\theta),1-\varepsilon,1+\varepsilon)\hat{A}_{t})]\end{split}$ (1)
where $w_{t}(\theta)=\frac{\pi_{\theta}(o_{t}\mid q,o_{<t})}{\pi_{\theta_{\text{old }}}(o_{t}\mid q,o_{<t})}$. Here, $(q,a)$ denotes question-answer pairs from dataset $\mathcal{D}$. $o_{t}$ and $o_{<t}$ represent the response token at position $t$ and all response tokens before $t$. The current and previous policies are parameterized by $\pi_{\theta}$ and $\pi_{\theta_{\text{old}}}$ respectively. The advantage estimator $\hat{A}_{t}$ employs Generalized Advantage Estimation (GAE) (Schulman et al., 2015) with an outcome-based reward function $R$ and a value function $V$, while $\varepsilon$ controls the clipping range. By treating question-response prefixes $q,o_{<t}]$ as distinct states and constraining policy updates within a trust region, PPO stabilizes RL training for LLMs. However, the joint optimization of policy and value functions introduces computational overhead, limiting training efficiency.
2.1.2 Group Relative Policy Optimization (GRPO)
@@
$\begin{split}\mathcal{J}_{\mathrm{GRPO}}(\theta)=\mathbb{E}_{(q,a)\sim\mathcal{D},\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old }}(\cdot\mid q)}}[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}(\\ \min(w_{i,t}(\theta)\hat{A}_{i},\operatorname{clip}(w_{i,t}(\theta),1-\varepsilon,1+\varepsilon)\hat{A}_{i})\\ -\beta D_{\mathrm{KL}}(\pi_{\theta}\|\pi_{\mathrm{ref}}))].\end{split}$ (2)
where $w_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}{\pi_{\theta_{\text{old }}}(o_{i,t}\mid q,o_{i,<t})}$ and $\hat{A}_{i}=\frac{R_{i}-\operatorname{mean}(\{R_{i}\}_{i=1}^{G})}{\operatorname{std}(\{R_{i}\}_{i=1}^{G})}$. $G$ responses $\{o_{i}\}_{i=1}^{G}$ are generated per input, $\pi_{\mathrm{ref}}$ denotes the reference policy, and $\beta$ controls KL penalty strength. GRPO replaces value function estimation with group-wise normalization on outcome-based rewards, improving training efficiency and effects through reduced computational complexity and group-wise sampling process. However, in most reasoning settings, the outcome reward is accuracy-based (Guo et al., 2025); for hard samples, most rollouts are incorrect and receive 0 (Yu et al., 2025), which provides limited training guidance within a group. As training progresses, these methods will quickly converge to similar response sequences and no longer provide novel answers, which may also lead to suboptimal performance (Xie et al., 2025).
@@
2.2 IMAGINE
As previously stated, despite advancements, PPO and GRPO face two key limitations: (1) Sparse reward signals that provide limited training guidance, and (2) weak exploration mechanisms during policy updates. These issues hinder learning efficiency and final optimization. To address these challenges, we propose an Intrinsic MotivAtion Guided exploratIoN for Enhanced reasoning (IMAGINE) method, which introduces structured exploration to enhance LLM reasoning and can be easily implemented in SOTA RL methods like PPO and GRPO. As shown in Figure 1, vanilla RL methods with outcome-based rewards focus only on correct trajectories, ignoring under-visited ones and causing LLMs to converge suboptimally with sparse rewards. IMAGINE additionally rewards incorrect under-visited trajectories, preventing the model from getting stuck via intrinsic motivation guided exploration.
Figure 3: Response length comparison on Countdown-4.
A Case Study is presented in Figure 2 to illustrate the significance of IMAGINE. Specifically, vanilla GRPO-trained LLM exhibits logical errors (e.g., missing numbers, redundancy in ①), repetitive reasoning patterns (②) due to intensive exploitations without explorations, and failures to solve tasks (③) or apply critical operations like multiplication (④). With the guidance of the exploration rewards from IMAGINE, LLM trained with IMAGINE (GRPO) reduces logical errors, diversifies reasoning paths (④), and results in a correct answer (③), demonstrating improved reasoning ability. Additionally, as shown in Figure 3, IMAGINE increases the average response length during training, stabilizing at a higher level than vanilla RL methods. This suggests that IMAGINE ’s exploration strategies enable LLMs to learn more complex reasoning processes with longer CoTs through considering diverse reasoning paths in responses, leading to performance gains. In the following subsections, we will systematically introduce IMAGINE through detailing its key components.
@@ 2.2.1 Trajectory-aware Exploration Rewards
IMAGINE provides intrinsic exploration rewards for LLM reasoning using a simple novelty-by-prediction mechanism. This approach is computationally efficient and does not require task-specific prior knowledge. It maintains two lightweight networks: a randomly initialized and frozen target network $f_{\theta_{T}}$ and a trainable predictor $f_{\theta_{P}}$ (with the same architecture). The predictor is trained to match the target on sequences observed during training:
$\mathcal{L}(\boldsymbol{s})=\|f_{\theta_{P}}(\boldsymbol{s})-f_{\theta_{T}}(\boldsymbol{s})\|_{2}^{2}$ (3)
where $\boldsymbol{s}$ denotes the current “state” in the RL paradigm. Since the two networks share the same architecture, this loss is guaranteed to converge (Burda et al., 2018). In the autoregressive LLM generation setting, it corresponds to the current generated question-response prefixes $q,o_{\leq t}]$. Intuitively, if a sequence has been visited frequently, the predictor can fit the target well, and the error becomes small; if it is rarely visited, the error remains large. Therefore, as illustrated in previous studies (Burda et al., 2018), the prediction error could be applied as an intrinsic novelty signal:
$R^{nov}(\boldsymbol{s})=\frac{\|f_{\theta_{P}}(\boldsymbol{s})-f_{\theta_{T}}(\boldsymbol{s})\|_{2}^{2}}{\operatorname{std}(\|f_{\theta_{P}}(\boldsymbol{s})-f_{\theta_{T}}(\boldsymbol{s})\|_{2}^{2})}.$ (4)
@@
The denominator uses a moving estimate of the standard deviation during training to keep the reward scale stable, and remains fixed during evaluation for determinism.
Unresolved Challenges: However, through careful empirical analysis, we identify that naively applying token-level novelty signals to LLMs results in severe performance degradation and computational inefficiency. In autoregressive LLM reasoning, when states are defined at the token level (i.e., $\boldsymbol{s}_{t}=[q,o_{\leq t}]$), which represents the standard paradigm for autoregressive generation, a single sampled response $o$ would produce a total number of $|o|$ states $\{\boldsymbol{s}_{t}\}_{t=1}^{|o|}$ of varying lengths for predictor update. This formulation introduces two critical practical challenges: (1) Dynamic episodic length. The distribution of prefix lengths exhibits systematic bias: shorter prefixes occur in every sampled response, whereas longer prefixes appear exclusively in extended responses. This leads to severe under-sampling of longer sequences, as illustrated in Figure 4. (2) Computational overload. Token-level state processing necessitates $O(|o|)$ forward passes and predictor updates per sampled response, resulting in prohibitive computational costs that scale quadratically with the number of rollouts.
These phenomena are evident in Figure 4: for 5 rollout responses of one question with lengths $[253,416,515,549,701]$, token-level state processing already yields $253+416+515+549+701=2434$ inputs for predictor update, dominated by short prefixes.
To address these challenges, we propose trajectory-aware exploration rewards operating at the sequence level. Specifically, we keep the same dual-network design but process the complete response sequence $(q,o)$ as a single input instead of using each $q,o_{\leq t}]$:
@@
$\left\{\begin{aligned} &\mathcal{L}(q,o)=\|f_{\theta_{P}}(q,o)-f_{\theta_{T}}(q,o)\|_{2}^{2}\\ &R^{\star}_{1}(q,o)=\frac{\|f_{\theta_{P}}(q,o)-f_{\theta_{T}}(q,o)\|_{2}^{2}}{\operatorname{std}(\|f_{\theta_{P}}(q,o)-f_{\theta_{T}}(q,o)\|_{2}^{2})}\end{aligned}\right$. (5)
@@
Here, $\mathcal{L}(q,o)$ is used to update $f_{\theta_{P}}$, while $R^{\star}_{1}(q,o)$ provides a single exploration reward for each complete response. This design eliminates token-level bias through uniform sequence treatment and reduces computation from $O(|o|)$ to $O(1)$ per training sequence, while remaining consistent with RLVR outcome-based reward paradigms.
Figure 4: Input statistics for predictor update.